# The advent of fully variational quantum eigensolvers using a hybrid multiresolution approach

### v1

@@ Abstract
In electronic structure theory, variational computing offers a valuable paradigm for the approximation of electronic ground states. However, for historical reasons, this principle is mostly restricted to model-chemistries in pre-defined fixed basis sets. Especially in quantum computation, these model-chemistries are far from an accurate description of the initial electronic Hamiltonian. Throwing down the gauntlet, we here demonstrate a fully variational approach to the electronic structure problem by variationally optimizing the orbitals representing the second quantized Hamiltonian alongside a quantum circuit that generates the many-electron wavefunction. To this end, the orbitals are represented within an adaptive multi-wavelet format, guaranteeing numerical precision. We then showcase explicit numerical protocols and highlight the quantum circuit’s effects on determining the optimal orbital basis.
@@ II Fully-Variational Quantum Chemistry
Since the orbital update of each orbital also depends on all other orbitals, the update is performed iteratively until all orbitals converge. For faster convergence, more efficient solvers can be used instead of the fixed point update shown in Eq. (23) – historically the KAIN [71] approximation has proven useful here. Since the update of all orbitals is performed simultaneously without updating the integrals within an iteration step, the orthonormality of the orbitals is no longer given after an iteration step, which makes a reorthonormalization after each iteration step necessary. To this end, we make use of the symmetric (“Löwdin") orthogonalization following Ref. 25. After convergence of all orbitals, the new and tensors are used to update the Hamiltonian of the quantum part for optimization of the many-body wavefunction ). The procedure of adapting the many-body wavefunction and orbital refinement is then repeated sequentially until the energy converges to a given threshold. We illustrate our proposed multiresolution quantum computing framework in the flowchart 1.
@@ III Numerical Demonstrations
We commence by highlighting basic features of the orbital refinement compared to the previous PNO approximation on simple (effective) two-electron model systems (H2, LiH) before continuing with chemically more interesting systems (BeH2 and H4). In the latter, we will in particular highlight the influence of an approximate VQE Ansatz (leading to a qubit wavefunction with varying quality across the molecular instances) on the basis refinement.
@@
III.1 H2 and LiH

### v2

@@ Abstract
In electronic structure theory, variational computing offers a valuable paradigm for the approximation of electronic ground states. However, for historical reasons, this principle is mostly restricted to model-chemistries in pre-defined fixed basis sets. Especially in quantum computation, these model-chemistries are far from an accurate description of the initial electronic Hamiltonian. This work demonstrates a fully variational approach to the electronic structure problem by variationally optimizing the orbitals representing the second quantized Hamiltonian alongside a quantum circuit that generates the many-electron wavefunction. To this end, the orbitals are represented within an adaptive multi-wavelet format, guaranteeing numerical precision. We then showcase explicit numerical protocols and highlight the quantum circuit’s effects on determining the optimal orbital basis.
@@ I Technical Background
This work builds up from a series of prior works that all focus focusing on the determination of numerically exact real-space representations of chemical systems in a black-box fashion. Further below, the broader background is illuminated, while we summarize the developments most important for this work in this paragraph.
In Ref. [11] Kottmann, Bischoff and Valeev realized the first attempt to determine optimal orbitals in an adaptive real-space basis from scratch, essentially reaching the basis set limit with affordable runtime. Here the correlated many-body method was Møller-Plesset Perturbation Theory (MP2) and the determined orbitals where pair-natural orbitals (PNOs). As a follow-up, Kottmann, Aspuru-Guzik et.al. employed the optimal PNOs to construct second-quantized Hamiltonians that can then be tackled with the full toolset of classical quantum chemistry (e.g. coupled-cluster or configuration interaction) as well as variational quantum algorithms, realizing the first hybrid approach that couples adaptive real-space orbital determination with general many-body methods. [12] Here, the notion of a surrogate functional, to determine the optimal orbital-basis was introduced – in this methodology, the basis error is solely determined by the quality of the surrogate functional. In a further work, the same authors adapted the the variational quantum circuits (many-body part) to the surrogate model (one-body-part) leading to the separable pair-approximation. [13]
In Ref. [14] Valeev,Harrison,et. al. derived the equations to determine optimal orbitals for general many-body methods. In the terminology of the previous works these equation determine an exact surrogate potential depending to the reduced density matrices of an arbitrary many-body method.
This work functions as a bridge between the surrogate functional of [12] and Valeev’s functional [14] where we focus on the application domain of quantum computation and interdependence of the quality of the determined orbital-bases with the used quantum circuits.
After initial release of this work, Nibbi, Frediani, Dinvay and Mendl released a preprint [15] investigating the combination of Valeev’s functional with DMRG solvers for the many-body part.
We note that different approaches formulated in real-space [16, 17, 18], as well as methods to directly optimize Gaussian based basis-sets (e.g. pioneering works of Tamayo-Mendoza et. al. [19] using automatic differentiation, see also [20, 21] and recent works [22]) have been investigated as well.
In the following we will give an overview over the individual fields of research essential to this work.
@@ II Fully-Variational Quantum Chemistry
In a general form, Eq. (23) can be written as
with a potential operator for each orbitals that is also functional of all orbitals. Using the terminology of Ref. [12] this is the surrogate potential that determines the basis of the electronic system. Valeevs potential (23) denotes an exact potential, while the PNO-based potential of Ref. [11] are an approximation. In the same way, a mean-field potential could be used, but this typically results in unbounded virtual orbitals – i.e. for most molecular systems we would only get bound solutions.
Since the orbital update of each orbital also depends on all other orbitals, the update is performed iteratively until all orbitals converge. For faster convergence, more efficient solvers can be used instead of the fixed point update shown in Eq. (23) – historically the KAIN [76] approximation has proven useful here. Since the update of all orbitals is performed simultaneously without updating the integrals within an iteration step, the orthonormality of the orbitals is no longer given after an iteration step, which makes a reorthonormalization after each iteration step necessary. To this end, we make use of the symmetric (“Löwdin") orthogonalization following Ref. 13. After convergence of all orbitals, the new and tensors are used to update the Hamiltonian of the quantum part for optimization of the many-body wavefunction ). The procedure of adapting the many-body wavefunction and orbital refinement is then repeated sequentially until the energy converges to a given threshold. We illustrate our proposed multiresolution quantum computing framework in the flowchart 1.
@@ III Numerical Demonstrations
We commence with one of the largest possible systems for the orbital refinement implementation presented in this work, cyclohexane with 36 electrons active electrons in 36 spatial (72 spin-) orbitals. Subsequently, we highlight the fundamental features of orbital refinement compared to the previous PNO approximation on simple (effective) two-electron model systems (H2, LiH) before continuing with chemically more interesting systems (BeH2 and H4). In the latter, we will in particular highlight the influence of an approximate VQE Ansatz (leading to a qubit wavefunction with varying quality across the molecular instances) on the basis refinement.
@@
III.1 Cyclohexane
We used the cyclohexane molecule as a demonstration of a molecule that is still feasible to compute with exact basis refinement and will give an exemplatory illustration of the cost and perspective gains of orbital refinement. The results are shown in Fig. 2 To the best of our knowledge, this is one of the largest simulations for VQE optimizations as well as exact orbital determination.
When compared to the cheaper MRA-PNOs, the total error is between 20-40 mEH depending on the used many-body state: in this case either a SPA quantum cirucit or a DMRG calculation with large bond dimension. The determination of the MRA-PNOs takes around 15 minutes on a standard compute node, while the full refinement coupled to a DMRG calculation takes around 1-2 days. These timings should not be taken as accurate benchmarks but rather as illustrations of the necessary cost to reach the last 20-40 mEh in the energy. In this case, a full basis refinement with quantum circuits on SPA level alone is most likely not an attractive scenario, as the method-error of the SPA circuits alone is already around 100 mEh which is significantly higher as the corresponding basis error of around 20 mEh.
To put the numbers into perspective: An orbital-optimized active space simulation within the cc-pVDZ basis has already a significantly larger basis error of around 100 mEh (75 of them already at the mean-field level). In our case the calculation took around 15 hours and was significantly longer than the MRA-PNOs that took around 15 minutes (with around the same time for the subsequent many-body simulation).
III.2 H2 and LiH
@@ Appendix A Computational Details and Notation
![[Uncaptioned image]](x2.png)
![[Uncaptioned image]](x3.png)
![[Uncaptioned image]](x4.png)
![[Uncaptioned image]](x5.png)
![[Uncaptioned image]](x12.png)
![[Uncaptioned image]](x13.png)
![[Uncaptioned image]](x14.png)
![[Uncaptioned image]](x15.png)